OpenStudy (anonymous):

On a stretch ofsingle-lane road with no entrances or exits the traffic density r(x,t)is a continuous function of distance xand time t, for all t > 0, and the traffic velocity u(r) is a function of density alone. Two alternative models are proposed to represent u: (i) u= uSL (1 – r^n/ r^nmax) where n is a positive constant (ii) u= 1/3 uSL In (rmax/r)where uSLis the maximum speed limit on the road and rmax is themaximum density of traffic possible on the road (corresponding tobumper-to-bumper traffic). a) Evaluate the maximum rate of traffic flow for cases (i)and (ii) above. Show t

OpenStudy (anonymous):

Consider a long, thin, uniform rod of constant cross-section whose temperature distribution is θ(x,t), surrounded by an atmosphere of constant uniform temperature θ0. The sides of the rod are not insulated so that heat is lost from the longitudinal surface at a rate h(θ(x,t)- θ0) per unit length, for some constant h. Derive the equation which describes the temperature distribution in the rod and show that it can be written ∂φ/∂t=k/ρc (∂^2 φ)/(∂x^2 ) - (h/ρAc) φ where φ(x,t) = (θ(x,t)- θ0), k is the thermal conductivity, ρ the density , c the thermal capacity and A the cross-sectional area of the rod. (You may assume one-dimensional heat flow in the rod.) Such a rod of length l is initially maintained at a constant temperature θ1. Both ends of the rod are then brought to and maintained at temperature θ0 and the rod is surrounded by an atmosphere at temperature θ0. Show that the temperature distribution in the rod at any subsequent time t has the form θ(x,t)= θ_0+∑_(n=1)^∞〖B_n e^(-k/ρc) (n^2 π^2)/l^2 +h/KA)t ) Sin (nπx/l)〗 and determine Bn. (When attempting to solve the equation above you may find it easier to swap the φ term over to the other side of the equation.)